OpenStudy (anonymous):

What is the solution of log((3x + 4))4096 = 4? ( the 3x+4 is like an exponent but lower

OpenStudy (anonymous):

write the equation (exponential) and then look...

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (anonymous):

\[\log \,_{(3x+4)}\,4096=4\Rightarrow(3x+4)^4=4096\]

OpenStudy (anonymous):

Yeah I didn't do this right at all, I got 1,280. I did 4^4=256 4096-256= 3,840/3

OpenStudy (anonymous):

? 8^4 = 4096 => 3x+4 = 8 (base must be positive)

OpenStudy (anonymous):

okay, i got the answer 4/3? but Where did you get the 8 from? o-o

OpenStudy (anonymous):

and thank you for your help by the way!

OpenStudy (anonymous):

(4096)^(1/4)